OpenStudy (anonymous):

If m = 7/15 and n = 3/2, then what is m/n?

OpenStudy (calculusxy):



hero (hero):

\[\dfrac{m}{n} = \dfrac{\dfrac{7}{15}}{\dfrac{3}{2}}\]

OpenStudy (calculusxy):

The way that Hero is using is means and extremes.

OpenStudy (calculusxy):